# KV-Edit: Training-Free Image Editing for Precise Background Preservation

### v2

@@ 1 Introduction
To overcome all the above limitations, we propose a new training-free method that preserves background consistency during editing. Instead of relying on regular attention modification or new inversion samplers for similar results, we implement KV cache in DiTs [39] to preserve the key-value pairs of background tokens during inversion and selectively reconstruct only the editing region. Our approach first employs a mask to distinguish between background and foreground regions and then inverts the image into noise space while caching KV values of background tokens at each timestep and attention layer. During the subsequent denoising process, only foreground tokens are processed, while their keys and values are concatenated with the cached background information. Effectively, we guide the generative model to maintain new content continuity with the background and keep the background content identical to the input. We call this approach KV-Edit.
@@ 3.4 KV-Edit
Building upon Eq. 8, achieving background-preserving foreground editing requires providing appropriate key-value pairs for the background. Our core insight is that background tokens’ keys and values reflect their deterministic path from image to noise. Therefore, we implement KV cache during the inversion process, as detailed in Algorithm 1. This approach records the keys and values at each timestep and block layer along the probability flow path, which are subsequently retrieved during denoising based on the corresponding timestep and layer index as shown in Algorithm 2. We term this complete pipeline “KV-Edit”, where “KV” represents the key-value caching mechanism employed within the DiT architecture, as illustrated in Fig. 2. Since the background tokens are preserved rather than regenerated, KV-Edit ensures perfect background consistency, effectively circumventing the three influencing factors discussed in Sec. 3.2.
@@ 3.5 Memory-Efficient Implementation
However, the inversion-free method may occasionally result in content retention artifacts as shown in Fig. 5 and FlowEdit [23]. Since our primary focus is investigating background preservation during editing, we leave further analysis of the inversion-free approach for future work.

### v3

@@ 1 Introduction
To overcome all the above limitations, we propose a new training-free method that preserves background consistency during editing. Instead of relying on regular attention modification or new inversion samplers for similar results, we implement KV cache in DiTs [39] to preserve the key-value pairs of background tokens during inversion and selectively reconstruct only the editing region. Our approach first employs a mask to decouple attention between background and foreground regions and then inverts the image into noise space while caching KV values of background tokens at each timestep and attention layer. During the subsequent denoising process, only foreground tokens are processed, while their keys and values are concatenated with the cached background information. Effectively, we guide the generative model to maintain new content continuity with the background and keep the background content identical to the input. We call this approach KV-Edit.
@@ 3.4 KV-Edit
Building upon Eq. 8, achieving background-preserving foreground editing requires providing appropriate key-value pairs for the background. Our core insight is that background tokens’ keys and values reflect their deterministic path from image to noise. Therefore, we implement KV cache during the inversion process, as detailed in Algorithm 1. This approach records the keys and values at each timestep and block layer along the probability flow path, which are subsequently used during denoising as shown in Algorithm 2. We term this complete pipeline “KV-Edit” as shown in Fig. 2 where “KV” means KV cache.
Unlike other attention injection methods [50, 4, 49], KV-Edit only reuses KV for background tokens while regenerating foreground tokens, without requiring specification of particular attention layers or timesteps. Rather than using the source image as injected information, we treat the deterministic background as context and the foreground as content to continue generating, analogous to KV cache in LLMs. Since the background tokens are preserved rather than regenerated, KV-Edit ensures perfect background consistency, effectively circumventing the three influencing factors discussed in Sec. 3.2.
@@ 3.5 Memory-Efficient Implementation
However, the inversion-free method may occasionally result in content retention artifacts as shown in Fig. 5 and FlowEdit [23]. Since our primary focus is investigating background preservation during editing, we leave more discussion about inversion-free in supplementary materials.
@@ 5 Conclusion
Appendix
In this supplementary material, we provide more details and findings. In Appendix A, we present additional experimental results and implementation details of our proposed KV-Edit. Appendix B provides further discussion and data regarding our inversion-free methodology. Appendix C details the design and execution of our user study. Moreover, In Appendix D, we discuss potential future directions and current limitations of our work.
Appendix A Implementation and More Experiments
Figure A: Additional editing results of KV-Edit. Our method demonstrates robust performance with strict background preservation and high image quality. Users can achieve creative designs by simply adjusting text prompts and masks according to their needs.
Implementation Details. Our code is built on Flux [1], with modifications to both double block and single block to incorporate KV cache through additional function parameters. Input masks are first downsampled using bilinear interpolation, then transformed from single-channel to 64-channel representations following the VAE in Flux [1]. In the feature space, the smallest pixel unit is 16 dimensions rather than the entire 64-dimensional token. Therefore, in addition to KV cache, we preserve the intermediate image features at each timestep to ensure fine-grained editing capabilities. In our experiment, inversion and denoising can be performed independently, allowing a single image to be inverted just once and then edited multiple times with different conditions, further enhancing the practicality of this workflow.
Experimental Results. Due to space constraints in the main paper, we only present results on the PIE-Bench [21]. Here, we provide additional examples demonstrating the effectiveness of our approach. To further showcase the flexibility of our method, Fig. A and Fig. B present various editing target applied to the same source image, without explicitly labeling the input masks because each case corresponds to a different mask. Fig. D illustrates the impact of steps and reinitialization strategy on the color changing tasks and inpainting tasks.
When changing colors, as the number of skip-steps decreases and reinitialization strategy is applied, the color information in the tokens is progressively disrupted, ultimately achieving successful results. In our experiments, the optimal number of steps to skip depends on image resolution and content, which can be adjusted based on specific needs and feedback. Unlike previous training-free methods, our approach even can be applied to inpainting tasks after employing reinitialization strategy, as demonstrated in the third row of Fig. D. The originally removed regions in inpainting tasks can be considered as black objects, thus requiring reinitialization strategy to eliminate pure black information and generate meaningful content. We plan to further extend our method to inpainting tasks in future work, as there are currently very few training-free methods available for this application.
Attention Scale When dealing with large masks (e.g., background changing tasks), our original method may produce discontinuous images including conflicting content, as illustrated in Fig. C. Stable-Flow [4] demonstrated that during image generation with DiT [39], image tokens primarily attend to their local neighborhood rather than globally across most layers and timesteps.
Consequently, although our approach treats the background as a condition to guide new content generation, large masks can introduce generation bias which ignore existing content and generate another objects. Based on this analysis, we propose a potential solution as shown in Fig. C. We directly increase the attention weights from masked regions to unmasked regions in the attention map (produced by query-key multiplication), effectively mitigating the bias impact. This attention scale mechanism enhances content coherence by strengthening the influence of preserved background on new content.
Figure B: Additional editing results of KV-Edit. Our method demonstrates robust performance with strict background preservation and high image quality. Users can achieve creative designs by simply adjusting text prompts and masks according to their needs.
Figure C: Implementation of attention scale. The scale can be adjusted to achieve optimal results.
Figure D: Additional ablation studies on two tasks. The first and second rows demonstrate the impact of timesteps and reinitialization strategy (RI) on color changing. The third row demonstrates the impact of timesteps and RI on the inpainting tasks.
Appendix B More Discussions on Inversion-Free
We implement inversion-free editing on Flux [1] based on the code provided by FlowEdit [23]. As noted in FlowEdit [23], adding random noise at each editing step may introduce artifacts, a phenomenon we also demonstrate in the main paper. In this section, we primarily explore the impact of inversion-free methods on memory consumption.
Algorithm A demonstrates the implementation of inversion-free KV-Edit, where “KV-inversion” and “KV-denoising” refer to single-step noise prediction with KV cache. KV cache is saved during a one-time inversion process and immediately utilized in the denoising process. The final vector can be directly added to the original image without first inversing it to noise. This strategy ensures that the space complexity of KV cache remains $O(1)$ along the time dimension. Moreover, resolution has a more significant impact on memory consumption as the number of image tokens grows at a rate of $O(n^{2})$.
We conducted experiments across various resolutions and time steps, reporting memory usage in Tab. A. When processing high-resolution images and more timesteps, personal computers struggle to accommodate the memory requirements. Nevertheless, we still recommend the inversion-based KV-Edit approach for several reasons:
  1. 1.
    Current inversion-free methods occasionally introduce artifacts.
  2. 2.
    Inversion-based KV-Edit enables multiple editing attempts after a single inversion, significantly improving usability and workflow efficiency.
  3. 3.
    Large generative models inherently require substantial GPU memory, which presents another challenge for personal computers. Therefore, we position inversion-based KV-Edit as a server-side technology.
timesteps $512\times 512$ $768\times 768$
Ours +Inf. Ours +Inf.
24 steps 16.2G 1.9G 65.8G 3.5G
28 steps 19.4G 1.9G 75.6G 3.5G
32 steps 22.1G 1.9G 86.5G 3.5G
Table A: Memory usage at different resolutions and timesteps. Our approach has a space complexity of $O(n)$ along the time dimension, while inversion-free methods achieve $O(1)$.
Algorithm A Simplified Inf. version KV-Edit
1:Input: $t_{i}$, real image $x^{src}_{0}$, foreground $z^{fg}_{t_{i}}$,foreground region $mask$, KV cache $C$
2:Output: Prediction vector $V^{fg}_{\theta t_{i}}$
3:$N_{t_{i}}\sim\mathcal{N}(0,1)$
4:$x^{src}_{t_{i}}=(1-t_{i})x^{src}_{t_{0}}+t_{i}N_{t_{i}}$
5:$V^{src}_{\theta t_{i}},C=\text{KV-Inverison}(x^{src}_{t_{i}},t_{i},C)$
6:$\widetilde{z}^{fg}_{t_{i}}=z^{fg}_{t_{i}}+mask\cdot(x^{src}_{t_{i}}-x^{src}_{0})$
7:$\widetilde{V}^{fg}_{\theta t_{i}},C=\text{KV-Denosing}(\widetilde{z}^{fg}_{t_{i}},t_{i},C)$
8:Return $V^{fg}_{\theta t_{i}}=\widetilde{V}^{fg}_{\theta t_{i}}-V^{src}_{\theta t_{i}}$
Appendix C User Study Details
We conduct our user study in a questionnaire format to collect user preferences for different methods. We observe that in most cases, users struggle to distinguish the background effects of training-based inpainting methods (e.g., FLUX-Fill [1] sometimes increases grayscale tones in images). Therefore, we allowed participants to select “equally good” regarding background quality.
Additionally, PIE-Bench [21] contains several challenging cases where all methods fail to complete the editing tasks satisfactorily. Consequently, we allow users to select “neither is good” for text alignment and overall satisfaction metrics, as illustrated in Fig. E.
We implement a single-blind mechanism where the corresponding method for each question is randomly sampled, ensuring fairness in the comparison. We collect over 2,000 comparison results and calculate our method’s win rate after excluding cases where both methods are rated equally.
Figure E: User study. We provide a sample where participants were presented with the original image, editing prompts, results from two different methods for comparison and four questions from four aspects.
Appendix D Limitations and Future Work
In this section, we outline the current challenges faced by our method and potential future improvements. While our approach effectively preserves background content, it struggles to maintain foreground details. As shown in Fig. D, when editing garment colors, clothing appearance features may be lost, such as the style, print or pleats.
Typically, during the generation process, early steps determine the object’s outline and color, with specific details and appearance emerging later. In the contrast, during inversion, customized object details are disrupted first and subsequently influenced by new content during denoising. This represents a common challenge in the inversion-denoising paradigm [16, 50, 12].
In future work, we could employ trainable tokens to preserve desired appearance information during inversion and inject it during denoising, still without fine-tuning of the base generative model. Furthermore, our method could be adapted to other modalities, such as video and audio editing, image inpainting tasks. We hope that “KV cache for editing” can be considered an inherent feature of the DiT [39] architecture.